OpenStudy (anonymous):

2a+4b=6c solve for "b"

OpenStudy (shamil98):

2(a+2b)=6c a + 2b = 3c 2b = 3c - a b = (3c-a)/2 you want to isolate b by undoing what was done to it

OpenStudy (anonymous):

Oksy,

OpenStudy (texaschic101):

2a + 4b = 6c (solving for b) subtract 2a from both sides 4b = 6c - 2a divide by 4 b = (6c - 2a)/4 b = 6c/4 - 2a/4 b = 3/2c - 1/2a

OpenStudy (texaschic101):

which is the same as : b = (3c - a)/2

OpenStudy (anonymous):

alrighty